# Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge

### v1

@@
Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge
@@
In this study, we delve into the hidden threats posed to text watermarking by users with cross-lingual knowledge. While most research focuses on watermarking methods for English, there is a significant gap in evaluating these methods in cross-lingual contexts. This oversight neglects critical adversary scenarios involving cross-lingual users, creating uncertainty regarding the effectiveness of cross-lingual watermarking. We assess four watermarking techniques across four linguistically rich languages, examining watermark resilience and text quality across various parameters and attacks. Our focus is on a realistic scenario featuring adversaries with cross-lingual expertise, evaluating the adequacy of current watermarking methods against such challenges.
Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge
@@ 1 Introduction
The advancement of Large Language Models (LLMs) has significantly transformed text generation across various domains, producing outputs that often closely mimic human writing. This advancement has raised concerns within the research community regarding potential misuses, including academic misconduct, the spread of disinformation, and the creation of synthetic training data (Xue et al., 2024b; Bender et al., 2021; Xue et al., 2024c; Zheng et al., 2024; Xue et al., 2024a; Cai et al., 2024). In response to these challenges, watermarking methods have been developed to differentiate between human-written and AI-generated texts (Aaronson, 2022; Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; He et al., 2024; Dathathri et al., 2024; Chang et al., 2024).
@@
Watermarking involves embedding a signal in AI-generated texts to identify the generating LLM using hypothesis testing. Specifically, watermarking offers theoretical guarantees regarding the detectability of the embedded signal by performing statistical inference on the generated text and testing against the null hypothesis. Since watermarking alters the generation of the original LLM, it is crucial to ensure that the impact on text quality is minimal while maintaining the watermark’s detectability. (1)(1)(1)Although steganography and watermarking share the practice of embedding signals, steganography primarily aims to conceal information through alterations of the text’s meaning, whereas watermarking serves to assert text source without changing the text’s semantics.
@@
Much of the existing literature on watermarking Zhang et al. (2025) has primarily focused on the quality and detectability of watermarked texts, with a predominant emphasis on English texts. While the theoretical principles of these methods can be applied across languages, cross-lingual studies can reveal new adversarial scenarios in which watermark signal could be removed, and that have yet to be thoroughly investigated.
@@
A recent work by He et al. (2024) addressed adversarial scenarios within a cross-lingual context, examining the consistency of watermarking through translation attacks, and developed defenses in this specific setting. Other existing studies on watermarking robustness have largely regarded back translation as the primary method for examining the interplay between languages, often overlooking other potential adversarial scenarios that may emerge following a translation to non-English languages  (Kuditipudi et al., 2023; Zhao et al., 2023; Pang et al., 2024). This oversight accounts for both unintentional alterations made by users for clarity or context adjustment and sophisticated adversaries intentionally refining the text to destroy watermark traces. In this paper, we evaluate four representative watermarking methods under two high-level themes. First, syntactical watermarking, which involves syntax changes to the generated text by manipulating the log-probabilities (logits) Lou et al. (2022) before the decode stage. Second, Semantic watermarking, which involves semantics manipulation of the generated text before the decode stage as well. The syntactical methods we consider are KGW  (Kirchenbauer et al., 2023a), EXP  (Aaronson, 2022), and Unigram  (Zhao et al., 2023). For the semantic methods, we choose XSIR  (He et al., 2024) as it aligns with our cross-lingual investigation in this paper. Specifically, this paper addresses the following research questions.
@@
    RQ3: How do different adversarial approaches to cross-lingual watermark evasion perform? Specifically, how do attacks fare under our novel asymmetrical threat model, where the output language differs from the pivotal language?
Our investigation reveals that current watermarking methods face significant challenges in cross-lingual settings. While methods like XSIR shows promise for cross-lingual scenarios, all current approaches demonstrate vulnerabilities to different translation-based attacks. Furthermore, we find that traditional quality metrics may not adequately capture the diversity of cross-lingual watermarked text, necessitating new evaluation approaches. The main contributions of this paper are as follows:
@@
    We investigate the role of text watermarking across four different languages and four SOTA text watermarking methods, with a concentration on analyzing watermark detectability and quality.
@@ 2 Background and Related Work
Syntactical-based Watermarking. Syntactical-based watermarking involves editing the generated text by manipulating the log-probabilities of a certain model. Multiple studies in the literature has investigated this type by introducing hypothesis testing. Pioneering in this area, the studies in KGW (Kirchenbauer et al., 2023a, b) and EXP (Aaronson, 2022) generate the watermark as a function of hashing the previous $k-1$-gram of a prompt to generate the next $k$ token. They introduce hypothesis testing to provide theoretical guarantees of watermark detectability. This is done by performing a statistical inference on the generated text, and calculating a score that is compared against a threshold with a very low False Positive Rate (FPR). Consequently, a barrage of similar watermark studies have been conducted to investigate the robustness and quality of KGW and EXP watermarks. In Unigram Zhao et al. (2023), the authors argue the robustness of the watermark to removal attacks could be mitigated by hashing a pre-determined key (uni-gram) that is used for all generations. Kuditipudi et al. (2023); Christ et al. (2024) introduced distortion-free watermarking in which the distribution of watermarked and unwatermarked text is the same. Similarly, Dathathri et al. (2024) introduces speculative sampling to generate watermarks at scale. Lu et al. (2024); Lee et al. (2023) investigated the role of token entropy to the watermark detectability and quality and proposes methods for generating and detecting the watermark.
Semantic Watermarking. Syntactical watermarking can still be compromised by paraphrasing and back-translation watermark removal attacks. In response to this, a number of studies have been conducted to evaluate watermarking methods in such scenarios. SIR (Liu et al., 2024) introduced semantic hashing of previous context instead of tokens to combat security and robustness attacks related to k-gram methods. Consequently, He et al. (2024) introduced XSIR, which is an extension of SIR with cross-lingual settings in which semantically cross-lingual words are clustered together. Unlike SIR, XSIR is able to mitigate translation attacks between different languages. Most recently, the work in  Chang et al. (2024) introduces the use of watermarking in a blackbox setting where users can semantically mark generated output of closed-source LLMs without accessing the log-probabilities for intellectual property reasons. Similarly, Hou et al. (2023, 2024) introduce sentence-level clustering to semantically mark the generated output.
Post-hoc Detection Approaches. In our evaluate, we evaluate proactive detection methods in which the watermark signal is embedded in the text during generation. However, multiple studies have investigated passive/discriminator methods for AI-generated text detection. Tian (2023); Mitchell et al. (2023); Gehrmann et al. (2019) use some statistical patterns in AI-generated text and trained discriminator models that differentiate human-written from AI-generated texts. Alshammari and Elleithy (2024) introduces a method to detect AI-generated text in Arabic language by leveraging language-specific diacritics. Abdelnabi and Fritz (2021) investigates adversarial text watermarking by modifying the inner workings of a transformer model to embed a hidden watermark.
@@ 3 Threat Model
In our threat model for text watermarking, we adopt a practical pipeline for potential watermark removal attacks, recognizing the various methods an adversary might employ. This pipeline includes stages such as simple translation, translation followed by paraphrasing, and a combination of translation, paraphrasing, and translating back again to the original language as seen in Figure 1. We consider a setting where LLMs are closed-source, in which case the model owners are the main victim of watermark removal attacks.
Attacker Knowledge. We consider a multilingual adversary who may operate at varying levels of sophistication. The adversary might be naive, unaware of the presence of watermarks, or advanced, intentionally attempting to evade detection. They understand how to utilize technological tools that can alter text, potentially impacting the embedded watermarks. It’s crucial to consider that regular users might unintentionally remove watermarks by prompting an LLM directly in one language before translating and making textual edits as can be seen in Figure 1 (b) at stage ②. Ignoring this setting can mislead focus in the literature toward only intentional attacks utilizing pivotal languages as shown in Figure 1 (a) stage ②.
Attacker Capability. The adversary is equipped with the ability to access AI chat interfaces, through which they begin prompting and engaging with content. They can leverage translation APIs to translate texts into different languages, which may weaken or eliminate watermark signals. Moreover, they have the capability to use probably less sophisticated LLM to paraphrase the text, thereby further disrupting the watermark signal.
This comprehensive threat model highlights the challenges that the watermarked text goes through in a cross-lingual setting. Our threat model underscores the necessity of developing sophisticated watermarking methods that can endure these intricate and pipelined attacks.
@@ 4 Methodology
1.4 Quality Evaluation
Evaluating watermarked text quality is essential for assessing watermark methods. Most studies focus on English, so we introduce a pipeline for multi-lingual LLMs. Initially, we used perplexity (PPL), but it wasn’t effective across languages. PPL results can be shown in Appendix A, Figures 5 and  6. For better assessment, we use GPT-Judger from Singh et al. (2023) with an advanced OpenAI model. Due to k-gram repetitions in watermarked text, we also apply Self-BLEU to measure diversity and repetition. We combine these methods to create a metric that adjusts diversity more effectively than relying solely on Self-BLEU.
GPT-Judger. Using an external judge to examine the quality of text has proven to be useful  (Singh and Zou, 2023; Zheng et al., 2023), especially when text quality metrics like perplexity are unable to capture all aspects of text quality. GPT-Judger covers a range of criteria that reflect the essential parts of text quality. These categories include
Relevance to the prompt, Depth of detail, Clarity of writing, Coherence and logical flow, Originality and insight, Use of specific examples, and Accuracy of information.
Additionally, from GPT-judger, we could extract soft wins and highlight which quality criterion of the text is affected the most by the watermarking method.
Self-BLEU SB. SB Zhu et al. (2018) measures the diversity in a generated text by measuring how similar sentences are to each other in the same text. SB is calculated by matching a sentence (hypothesis) with all other sentences (references) in the same text for each sentence in the text. Then an average score is calculated and returned as the final score.
Adjusted Diversity AD. Because SB can be misleading in measuring the diversity of text –for example mixing tokens from different languages yields very low SB score, hence more diversity in text– we create a new diversity metric that reflects unwanted diversity or excessive repetition in text by leveraging the Judger’s coherency criterion scores as follow:
@@ 2.4 Watermark Methods
KGW (Kirchenbauer et al., 2023a). The KGW method embeds a watermark signal in generated text by manipulating log-probabilities (logits) of next token. The vocabulary $v$ is divided into green and red lists based on a split ratio $\gamma$. A secret key $S_{k}$ and a hash of the previous $k-1$ token ids seed a pseudorandom generator to produce the next token $k$. The generating model is either limited to the green list (hard watermark) or biased toward it (soft watermark) by adding a small value $\delta$ to the logits. Soft watermarking manages low-entropy contexts with few green list options.
@@
EXP Aaronson (2022). EXP uses exponential minimum sampling, which is a variant of the Gumbel trick  Papandreou and Yuille (2011), to bias the distribution of the next token generation. Specifically, like KGW, EXP uses $k-1$ context-window to seed a pseudorandom generator (PRG) to generate the next token $k$. However, instead of using soft watermark or applying a $\delta$ value to the logits, EXP uses PRG to generate random numbers $r_{t,i}$, which is the same size as the vocabulary of the generating model. Then, at position $t$, token $i$ is sampled by maximizing the following quantity.
@@
when $p_{t,i}$ is very small, token $i$ will only be chosen if $r_{t,i}$ is close to one, which is very unlikely to happen. In terms of randomness, this sampling method will return the same token every time the same $k-1$ context is used for the PRG. The watermark is then detected by the following equation.
@@
where $E$ is a multilingual embedding model, and $\Delta$ is the function that determines the watermark logit bias distribution for all tokens in the vocabulary. By doing so, XSIR ensures that tokens predicted under similar semantic prefixes receive similar watermarking bias.
@@
2 : GPT-judge results. We compute the average of watermarked and unwatermarked scores for $500$ generations to find that the criterion the judge complains the most about was coherency decrease.
@@ 5 Experimental Setup
In all our experiments, we assess the strength of the watermark both before and after various attacks across four watermarking methods. Following previous work Kirchenbauer et al. (2023a); Zhao et al. (2023); Dathathri et al. (2024), we evaluate watermark detection using Receiver Operating Characteristic (ROC) curves, which illustrate the trade-off between the false positive rate (FPR) and false negative rate (FNR). FPR indicates that the generated text is falsely flagged as human written, which is more critical in this context. Although theoretical guarantees exist for the FPRs of the watermark methods we study, we adopt an empirical approach. Specifically, we automatically calculate a threshold based on the scores of unwatermarked text and then compare it against the scores of watermarked text. This method ensures that a fair and consistent threshold is applied, particularly in a multilingual setting. For further details, Appendix B reports additional performance metrics (TPR, TNR, FPR, FNR) derived from the thresholds used by the evaluated methods.
LLMs. We use several multi-billion parameter models to generate and evaluate watermarked texts. Due to the lack of availability of open-source LLMs that support all of the languages we used in our study, we used different models for different languages. For Chinese and Indonesian, we use the Sailor2 1B and 8B variants Sailor2 Team (2024) for text generation and perplexity evaluation(2)(2)(2)Perplexity results are provided in the appendices.. For Arabic, we generate text with the Jais family 6.7B model Inception (2024) and assess perplexity using Acegpt 7B Huang et al. (2023). English is generated on both models. We generate watermarked text in different hyper-parameter settings and different languages with an average of 10 GPU hours for KGW and Unigram, an average of 7 hours for XSIR, and 3 hours for EXP. We also use OpenAI GPT-4o-mini-2024-07-18 as the GPT-based judger which is slower but better than GPT-3.5-turbo. Each experiment took around 20 minutes for 500 generations to assess text quality with GPT-Judger. For translation attacks, we employ OPUS-MT models Tiedemann and Thottingal (2020); Tiedemann et al. (2022). To ensure consistency with watermark protocols (e.g., using the same generation vocabulary for detection), we generate English texts with both the Jais and Sailor2 models; translations (for Chinese, Indonesian, or Arabic) are derived from the original English versions generated by their respective LLMs.
Watermark Parameters. For KGW Kirchenbauer et al. (2023a) and Unigram Zhao et al. (2023) we use three green list ratios, $\gamma\in\{0.1,0.5,0.9\}$, and three bias values, $\delta\in\{2,5,10\}$. For KGW, we also use a lefthash with a context size of 1(3)(3)(3)Distinct from Unigram, where a key is hashed once and reused for subsequent generations.. When applicable, we follow previous studies by using a z-score threshold of $4$. For EXP Aaronson (2022), we set a context length of $4$ for the hashing and consider text watermarked if the $p$-value falls below $10^{-4}$. For XSIR He et al. (2024), approximately half the vocabulary is split into green and red lists, i.e., $\gamma=0.5$. Although the original XSIR uses a bias value of $\delta=1$, we vary the $\delta\in\{2,5,10\}$ to facilitate extensive evaluation; we also adopt their watermark threshold of $0.2$.
Unless stated otherwise, all experiments use $\gamma=0.5$ and $\delta=2.0$ for KGW, Unigram, and XSIR, as these values empirically provide the best balance between watermark strength and text quality. Additional results with varying parameters are presented in the appendices, which confirm with previous analyses.
@@
Dataset. For all experiments, we use $500$ examples from the C4 dataset Raffel et al. (2020), which is available in all evaluated languages. For English, we follow previous work by using the RealNewsLike split; for other languages, we use the main training split, as RealNewsLike is available only for English(4)(4)(4)https://huggingface.co.

### v2

# Evaluating the Robustness and Accuracy of Text Watermarking Under Real-World Cross-Lingual Manipulations

@@
Evaluating the Robustness and Accuracy of Text Watermarking Under Real-World Cross-Lingual Manipulations
@@
We present a study to benchmark representative watermarking methods in cross-lingual settings. The current literature mainly focuses on the evaluation of watermarking methods for the English language. However, the literature for evaluating watermarking in cross-lingual settings is scarce. This results in overlooking important adversary scenarios in which a cross-lingual adversary could be in, leading to a gray area of practicality over cross-lingual watermarking. In this paper, we evaluate four watermarking methods in four different and vocabulary rich languages. Our experiments investigate the quality of text under different watermarking procedure and the detectability of watermarks with practical translation attack scenarios. Specifically, we investigate practical scenarios that an adversary with cross-lingual knowledge could take, and evaluate whether current watermarking methods are suitable for such scenarios. Finally, from our findings, we draw key insights about watermarking in cross-lingual settings(1)(1)(1)code and data: https://github.com/SecureDL/xlingual_watermark_eval.
Evaluating the Robustness and Accuracy of Text Watermarking Under Real-World Cross-Lingual Manipulations
@@ 1 Introduction
The advancement of Large Language Models (LLMs) has significantly transformed text generation across various domains, producing outputs that closely mimic human writing. This advancement has raised concerns within the research community regarding potential misuses, including academic misconduct, the spread of disinformation, and the creation of synthetic training data (Bender et al., 2021; Xue and Lou, 2022; Zheng et al., 2024; Xue et al., 2023b; Lou et al., 2024). In response to these challenges, watermarking methods have been developed to differentiate between human-written and AI-generated texts  (Aaronson, 2022; Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; He et al., 2024; Dathathri et al., 2024; Chang et al., 2024).
@@
Watermarking involves embedding a signal in AI-generated texts to identify the generating LLM using hypothesis testing. Specifically, watermarking offers theoretical guarantees regarding the detectability of the embedded signal by performing statistical inference on the generated text and testing against the null hypothesis. Since watermarking alters the original LLM output, it is crucial to ensure that the impact on text quality is minimal while maintaining the watermark’s detectability. (2)(2)(2)Although steganography and watermarking share the practice of embedding signals, steganography primarily aims to conceal information through alterations of the text’s meaning, whereas watermarking serves to assert text source without changing the text’s semantics.
@@
Much of the existing literature on watermarking has primarily focused on the quality and detectability of watermarked texts, with a predominant emphasis on English texts. While these methods are theoretically language-agnostic, cross-lingual studies can reveal new adversarial scenarios in which watermark signal could be removed, and that have yet to be thoroughly investigated.
Existing studies on watermarking robustness have largely regarded the interplay between languages by back translation in which the English watermarked text is translated to some pivot language then by translating back to English, weakening the watermark signal  (Kuditipudi et al., 2023; Zhao et al., 2023; Pang et al., 2024; Ghanim et al., 2024). This overlooks other potential adversarial scenarios that may emerge following a translation to non-English languages.
A recent work by He et al. (2024) attempted to explore adversarial scenarios within a cross-lingual context with specific translation attacks (e.g., Cross-lingual Watermark Removal Attack CWRA) to bypass watermarking by first obtaining a response from an LLM in a pivot language, which is then translated into the target prompt language. As can be shown in Figure 1-a, we refer to this type of translation including back-translation as symmetrical, since an attacker does not utilize the pivot language. However, more practical adversarial scenarios are not studied, especially in the settings where users with cross-lingual knowledge might take to remove the watermark. As shown in Figure 1-b, we call this asymmetrical since the user may use other languages without intentionally removing the watermark. This oversight accounts for both unintentional alterations made by users for clarity or context adjustment and sophisticated adversaries intentionally refining the text to destroy watermark traces such as CWRA. Additionally, a broader assessment of text quality under watermarking techniques in diverse languages is scarce but necessary, particularly given linguistic differences.
@@
In this paper, we evaluate four representative watermarking methods under two high-level themes. First, syntactical watermarking, which involves syntax changes to the generated text by manipulating the logits before the decode stage. Second, Semantic watermarking, which involves semantics manipulation of the generated text before the decode stage as well. The syntactical methods we consider are KGW  (Kirchenbauer et al., 2023a), EXP  (Aaronson, 2022), and Unigram  (Zhao et al., 2023). For the semantic methods, we choose XSIR  (He et al., 2024) as it aligns with our cross-lingual investigation in this paper. Specifically, this paper addresses the following research questions.
@@
    RQ3: How do different adversarial approaches to cross-lingual watermark evasion perform? Specifically, how do attacks fare under our asymmetrical threat model, where the output language differs from the pivotal language?
Our investigation reveals that current watermarking methods face significant challenges in cross-lingual settings. Furthermore, we find that traditional text quality metrics may not adequately capture the diversity of cross-lingual watermarked text, necessitating new quality evaluation approaches. The main contributions of this paper are as follows:
@@
    We investigate the role of text watermarking across four different languages and four popular text watermarking methods, with a concentration on analyzing watermark detectability and quality.
@@ 2 Background and Related Work
Syntactical-based Watermarking. Syntactical-based watermarking manipulates log-probabilities of generated text. KGW (Kirchenbauer et al., 2023a, b) and EXP (Aaronson, 2022) pioneered this by generating watermarks based on hashing previous tokens, introducing hypothesis testing with low False Positive Rate (FPR) for detection. Subsequent research expanded on these approaches: Unigram (Zhao et al., 2023) used a pre-determined key for all generations;  Kuditipudi et al. (2023); Christ et al. (2024) developed distortion-free watermarking to enhance the quality of watermarked text;  Dathathri et al. (2024) introduced speculative sampling for generation at scale; and  Lu et al. (2024); Lee et al. (2023) explored token entropy’s role in watermark detectability.
Semantic Watermarking. To address paraphrasing and back-translation attacks that compromise syntactical watermarks, semantic approaches emerged. SIR (Liu et al., 2024) introduced semantic hashing of context rather than tokens. XSIR (He et al., 2024) extended this with cross-lingual settings to counter translation attacks.  Chang et al. (2024) developed a blackbox semantic watermarking for closed-source LLMs, while  Hou et al. (2023, 2024) introduced sentence-level clustering. However, their rejection sampling approach slows watermarked text generation compared to methods in our benchmark.
Post-hoc Detection Approaches. Our evaluation focuses on proactive detection with embedded watermarks, but passive/discriminator methods exist for AI-generated text detection.  Tian (2023); Mitchell et al. (2023); Gehrmann et al. (2019) use statistical patterns with discriminator models to differentiate human from AI written text.  Alshammari and Elleithy (2024) detects AI-generated Arabic text using diacritics, while  Abdelnabi and Fritz (2021) embeds hidden watermarks by modifying transformer internals.
Watermarking via Backdoors. Recent work has explored backdoor-based approaches for fingerprinting and watermarking LLM outputs. Xu et al. (2024) used a technique to embed model-specific fingerprints through backdoor triggers in instruction tuning. Unlike traditional watermarking that modifies token probabilities during inference Kirchenbauer et al. (2023a); Aaronson (2022); Kuditipudi et al. (2023), backdoor approaches implant distinctive generation patterns during model training (Qi et al., 2021; Kurita et al., 2020) or at inference time (Al Ghanim et al., 2023; Xue et al., 2023a; Zheng et al., 2023). This creates persistent model behaviors that can be detected with special inputs that trigger the backdoor.
@@ 3 Threat Model
We propose a threat model for text watermarking that addresses realistic removal scenarios. Our model includes various attack pipelines including translation, paraphrasing, and combinations thereof as illustrated in Figure 1. We focus on closed-source LLMs where model owners are the primary victims of watermark removal attacks.
Attacker Knowledge: We consider multilingual adversaries ranging from naive (unaware of watermarks) to advanced (deliberately trying to evade detection). Importantly, regular users may unintentionally remove watermarks through normal multilingual usage patterns.
Attacker Capability: Adversaries can access AI chat interfaces, translation APIs, and have the ability to paraphrase or edit text, potentially disrupting watermark signals.
Real-world Scenarios: In an academic setting, non-English speaking professors providing English questions while allowing answers in native languages, enabling students to use AI-generated content with translation. This is not uncommon especially for science subjects where English is the language of the content is in English. Another scenario could take place in media and information sharing in which the post/content is translated from English to local languages and edited for specific audiences, inadvertently removing watermarks. We call these types of translations as asymmetric as the language of the utilized generation is different from that of the prompt.
These scenarios are common yet rarely considered in current watermark threat models. Robust watermarking systems must address multilingual use cases rather than focusing solely on using a language as pivotal to remove the watermark without utilizing it.
@@ 4 Methodology
1.4 Benchmark Challenges
Quality Metrics. Evaluating watermarked text quality is essential for assessing watermark methods. Most studies focus on English, so we introduce a pipeline for multi-lingual LLMs. Initially, we used perplexity (PPL), but it wasn’t effective across languages. PPL results can be shown in Appendix A, Figures 5 and  6. For better assessment, we use GPT-Judger from Singh et al. (2023) with an advanced OpenAI model. However, since we assess watermarking in multiple languages, evaluating quality with one method may not be sufficient to capture any possible linguistics nuances not captured in previous monolingual works. For example, multiple studies have shown biases introduced by the so called llm-as-a-judge paradigm in which a LLM is given multiple options to choose from (Zheng et al., ; Pezeshkpour and Hruschka, 2023; Ye et al., 2024; Koo et al., 2023). Therefore, we augment our quality evaluation with an evaluation of the judger itself by conducting fairness studies for any possible biases toward a specific language. Additionally, due to k-gram repetitions in watermarking, we apply Self-BLEU (SB) (Zhu et al., 2018) to measure diversity and repetition. We combine the results from GPT-Judger and SB to create a metric that adjusts diversity more effectively, and provides insights into evaluating watermarked text quality in cross-lingual settings.
Text Interpolations in Cross-lingual Settings. Existing literature on watermark attacks has primarily focused on intentional attacks designed to remove watermarks. In these studies, non-English languages are typically used as pivotal languages for back-and-forth translation to weaken or remove the watermark signal. However, this approach overlooks more practical and nuanced cross-lingual usage scenarios. Watermarked text in multilingual settings faces challenges from translation and meaningful use in the intermediate languages. Adversaries with multilingual capabilities may utilize, edit, and distribute the content in various languages. These scenarios present more complex threats than simple pivotal translation attacks.
Adjusted Diversity AD. Because SB alone can be misleading in measuring the diversity of text –for example mixing tokens from different languages yields very low SB score, hence more diversity in text– we create a new diversity metric that reflects unwanted diversity or excessive repetition in text by leveraging the Judger’s coherency criterion scores as follow:
@@ 2.4 Watermark Methods
KGW (Kirchenbauer et al., 2023a). The KGW method embeds a watermark signal in generated text by manipulating log-probabilities (logits) of next token. The vocabulary $v$ is divided into green and red lists based on a split ratio $\gamma$. A secret key $S_{k}$ and a hash of the previous k1k-1 token ids seed a pseudorandom generator to produce the next token $k$. The generating model is either limited to the green list (hard watermark) or biased toward it (soft watermark) by adding a small value $\delta$ to the logits. Soft watermarking manages low-entropy contexts with few green list options.
@@
EXP Aaronson (2022). EXP uses exponential minimum sampling, which is a variant of the Gumbel trick  Papandreou and Yuille (2011), to bias the distribution of the next token generation. Specifically, like KGW, EXP uses k1k-1 context-window to seed a pseudorandom generator (PRG) to generate the next token $k$. However, instead of using soft watermark or applying a $\delta$ value to the logits, EXP uses PRG to generate random numbers $r_{t,i}$, which is the same size as the vocabulary of the generating model. Then, at position $t$, token $i$ is sampled by maximizing the following quantity.
@@
when $p_{t,i}$ is very small, token $i$ will only be chosen if $r_{t,i}$ is close to one, which is very unlikely to happen. In terms of randomness, this sampling method will return the same token every time the same k1k-1 context is used for the PRG. The watermark is then detected by the following equation.
@@
where $E$ is a multilingual embedding model, and $\Delta$ is the function that determines the watermark logit bias distribution for all tokens in the vocabulary. XSIR uses semantic clustering to assign consistent biases, ensuring tokens with similar prefixes or context receive similar watermarking bias.
@@
2 : GPT-judge coherency criterion results. We compute the average of watermarked and unwatermarked scores for $500$ generations. We use Jais model for Arabic and Sail for Chinese and Indonesian. We use Cohere to represent all languages.
@@ 5 Experimental Setup
Choice of Models. Due to scarcity of open-source LLMs that universally support our targeted languages, we employ different models suitable for different languages. For Chinese and Indonesian, we utilize the Sailor2 1B and 8B variants Sailor2 Team (2024). For Arabic text generation, we use the Jais family 6.7B model Inception (2024), while perplexity is assessed using Acegpt 7B Huang et al. (2023). For English, we generate text using both aforementioned models. The use of varying models across languages is essential to accommodate the lack of universally compatible models but ensures credible cross-lingual watermarking assessments. Recently, a model that support all our targeted languages was released  (Cohere et al., 2025) in which all of our targeted languages are supported(3)(3)(3)https://huggingface.co/CohereLabs/c4ai-command-r7b-12-2024. We use this model to add experiments for an instruction-following and text-completion task with more languages.
@@
Dataset Selection. For all experiments, we use $500$ examples from the C4 dataset Raffel et al. (2020), which is available in all evaluated languages. For English, we follow previous work by using the RealNewsLike split; for other languages, we use the main training split, as RealNewsLike is available only for English(4)(4)(4)https://huggingface.co. We also utilized a cleaned version of LFQA dataset from this work  Krishna et al. (2023). We use this dataset along with instruction-following prompting for information entropy analysis in different languages to further assess the quality of text with watermarking.
Watermarking Parameters. We evaluate watermarking across four methods: KGW, Unigram, XSIR, and EXP. For KGW and Unigram, parameters include green list ratios $\gamma\in{0.1,0.5,0.9}$ and bias values $\delta\in{2,5,10}$. A context size of 1 is applied for KGW’s lefthash. For EXP, a context length of 4 and a $p$-value threshold of $10^{-4}$ are used. XSIR divides the vocabulary with $\gamma=0.5$ and bias values $\delta\in{2,5,10}$. Standard settings of $\gamma=0.5$ and $\delta=2.0$ are typically used unless otherwise noted, based on optimal empirical results. Additional parameter results are detailed in the appendices.
Watermarking Metrics Following prior studies Kirchenbauer et al. (2023a); Zhao et al. (2023); Dathathri et al. (2024), watermark detection is analyzed via ROC curves, focusing on the trade-off between the false and true positive/negative rates. An empirical threshold is determined using a comparison between unwatermarked and watermarked text scores to enhance multilingual watermark consistency. See Appendix B for detailed performance metrics using specific threshold values.
Language Bias Assessment. We conducted controlled experiments to identify potential language biases in LLM judges. Starting with English examples from the C4 dataset, we used GPT-3.5-Turbo to create perfect translations in seven languages: English, Arabic, Chinese, Indonesian, Persian (Farsi), German, and Japanese. This approach established ground truth by ensuring identical content across all languages, allowing us to isolate language-specific biases. For methodological rigor, we randomized text presentation order and conducted multiple runs with different seeds. See Appendix C for complete experimental details.
Execution Details. Experiments involve generating watermarked text in various languages, with processes taking approximately 10 GPU hours for KGW and Unigram, 7 hours for XSIR, and 3 hours for EXP. For qualitative assessment, the GPT-4o-mini-2024-07-18 serves as the GPT-based judger, conducting quality evaluations in roughly 20 minutes for 500 generations. Since translations are heavily used in this work, we opt for the open-sourced and easy to use OPUS-MT models Tiedemann and Thottingal (2020); Tiedemann et al. (2022) which are used by previous works such as Kuditipudi et al. (2023).